### v1

@@ 4.1 State space model
Let us assume that the signal of interest is modelled as a continuous-time Itô process $\mathtt{X}(t)$ in $(\Omega,{\mathcal{F}},\mathbb{P})$, described by the stochastic differential equations (SDEs)
@@
with initial condition $X_{n-1}=\mathtt{X}(t_{n-1})=x^{\prime}$. For example, a plain Euler-Maruyama scheme yields
@@
The time grid (22) typically corresponds to the time instants where observations are collected and a likelihood function $g_{n}$ is available. We assume that the observations have the form
@@ 4.2 Constrained kernels
Let ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ be a constraint as described in Section 3.1. Sampling the constrained kernel $\hat{K}_{n}(x^{\prime},dx)$ in (8) is equivalent to simulating paths of the Itô process $\mathtt{X}(t)$ that evolves from $\mathtt{X}(t_{n-1})=x^{\prime}$ up to a point in the set ${\mathcal{C}}_{n}$ at time $t_{n}$, i.e., we simulate $\mathtt{X}(t_{n})$ conditional on the initial condition $\mathtt{X}(t_{n-1})=x^{\prime}$ and the terminal condition $\mathtt{X}(t_{n})\in{\mathcal{C}}_{n}$.
One straightforward approach is to simulate trajectories of (21) and then reject the samples $\mathtt{X}(t_{n})\notin{\mathcal{C}}_{n}$. However this can be extremely inefficient as the dimension $d_{x}$ increases. An alternative approach is to “guide” the Itô process to generate samples contained in ${\mathcal{C}}_{n}$ at time $t_{n}$. It is well known that the Doob $h$-transform [Rogers00]
@@
(24) ${\sf d}\mathtt{X}^{n}(t)=\left[a(\mathtt{X}^{n}(t),t)+\sigma^{2}(\mathtt{X}^{n}(t),t)\nabla_{x}\log h(\mathtt{X}^{n}(t),t)\right]{\sf d}t+\sigma(\mathtt{X}^{n}(t),t){\sf d}\mathtt{W}(t),\quad\mathtt{X}^{n}(t_{n-1})=x^{\prime},$
describes a modified Itô process $\mathtt{X}^{n}(t)$, with $t\in[t_{n-1},t_{n}]$, such that the law of $\mathtt{X}^{n}(t)$ coincides with the law of $\mathtt{X}(t)$ conditional on $\mathtt{X}(t_{n-1})=x^{\prime}$ and $\mathtt{X}(t_{n})\in{\mathcal{C}}_{n}$. Note, however, that $h(X,t)$ is intractable except for a few particular cases and it is also expensive to approximate.
@@
be a flow of superlevel sets of the Doob $h$-transform generated by some function $r:[t_{n-1},t_{n}]\mapsto(0,\infty)$, such that $\lim_{t\to t_{n}}r(t)=\infty$ in order to ensure $S_{n}(t_{n})={\mathcal{C}}_{n}$. Given $r(t)$, there is some $\epsilon_{r}\geq 0$ such that
@@
Assuming $\epsilon_{r}<<1$, the deterministic flow $S_{n}(t)$ captures most of the probability mass of the constrained process $\mathtt{X}^{n}(t)$ in Eq. (24). Intuitively, we can interpret it as a “high probability tube” connecting the initial condition $\mathtt{X}(t_{n-1})=x^{\prime}$ with the terminal set ${\mathcal{C}}_{n}$ specified by the constraint $\mathcal{C}$}.
@@ 4.3 Barrier-constrained kernels
While approximating the Doob $h$-transform $h(x,t)$ is possible (several recent papers [Chopin23, Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25] have tackled this computation, e.g., using local linearisation techniques or neural networks), we explore a simpler alternative where we construct a modified SDE that mimics the structure of (24) but replaces $\log h(x,t)$ by a simple-to-evaluate, differentiable “barrier function” $\log b(x,t)$. This barrier is designed to
@@
(25) ${\sf d}\mathtt{X}^{n}(t)=\left[a(\mathtt{X}^{n}(t),t)-\mu\sigma^{2}(\mathtt{X}^{n}(t),t)\nabla_{x}\log b(\mathtt{X}^{n}(t),t)\right]{\sf d}t+\sigma(\mathtt{X}^{n}(t),t){\sf d}\mathtt{W}(t),\quad\mathtt{X}^{n}(t_{n-1})=x^{\prime},$
where $\mu>0$ is a parameter that controls the strength of the penalization when $\mathtt{X}^{n}(t)\notin\hat{S}_{n}(t)$.
Let ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:\log g_{n}(x)\geq\log\upsilon_{n}\}$ be a superlevel set of the likelihood $g_{n}$, in agreement with the assumptions of Theorem 3.6. This construction implies that the sets $\hat{S}_{n}(t)$ depend on the log-likelihood $g_{n}$, hence it is useful to make the barrier depend on this quantity as well. This is relatively simple and can be done in a number of ways. In this paper we use a shifted soft-plus function with parameters $\varrho$ and $\kappa\geq 1$,
@@
Note that, for large $\kappa$,
$u_{\varrho,\kappa}(z)\approx\left\{\begin{array}[]{ll}0&\text{when}~~z<\varrho\\ z-\varrho&\text{when}~~z>\varrho\\ \end{array}\right.$.
The barrier function is then constructed as
@@ 5.1 Stochastic Lorenz 96 model
where $U_{n}\sim{\mathcal{N}}(0,\sigma_{y}^{2}I_{d_{y}})$ and $H$ is a $d_{x}\times d_{y}$ matrix and $X_{n}=\mathtt{X}(t_{n})$. For each independent simulation trial, we generate $H$ randomly as
@@ 5.2 Barrier function
where $Y_{n}=y_{n}$ is the observation at time $t_{n}=n\Delta_{o}$. The constraint consists of superlevel sets ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:\log g_{n}(x)\geq\log\upsilon\}$. The tube $\hat{S}_{n}(t)$ is constructed as
$S_{n}(t):=\{x\in{\mathcal{X}}:\log g^{n}(x,t)\geq\log\upsilon\},\quad t\in[t_{n-1},t_{n}],$
@@
Then, by construction, $g^{n}(x,t_{n-1})=g_{n-1}(x)$, $g^{n}(x,t_{n})=g_{n}(x)$, $S_{n}(t_{n-1})={\mathcal{C}}_{n-1}$ and $\hat{S}_{n}(t_{n})={\mathcal{C}}_{n}$.
@@
$\hat{S}_{n}(t)=\{x\in{\mathcal{X}}:\log g^{n}(x,t)\geq\log\upsilon\}=\left\{x\in{\mathcal{X}}:\|y^{n}(t)-H^{\top}x\|\leq\sqrt{-\frac{\log\upsilon}{2\sigma_{y}^{2}}}\right\},$
hence, $\hat{S}_{n}(t)$ is the preimage (on the state space $\mathcal{X}$}) of the ball $\{y\in{\mathcal{Y}}:\|y^{n}(t)-y\|\leq R\}$, with radius $R=\sqrt{-\frac{\log\upsilon}{2\sigma_{y}^{2}}}$ (on the observation space $\mathcal{Y}$}). The barrier function in (26) becomes
@@
where $\varrho=-\log\upsilon$,
@@ 5.3 Stability of constrained filters
From the SSMs ${\mathcal{S}}^{A}$ and ${\mathcal{S}}^{B}$ we construct constrained models $\hat{\mathcal{S}}^{A}$ and $\hat{\mathcal{S}}^{B}$. The constrained kernels $\hat{K}_{n}$ are implemented using the barrier method of Section 4.3 with parameters $\varrho=2$, $\kappa=1$ and $\mu=10$. We simulate the Lorenz 96 model with dimension $d_{x}=10$, with $d_{y}=6$ observations at every time $t_{n}$, and noise scale parameters $\sigma_{x}^{2}=\frac{1}{2}$ and $\sigma_{y}^{2}=\frac{1}{4}$. The discretisation of the TV distance is carried out with $r=6$ and $r_{\rm sub}=3$. The filters $\hat{\pi}_{n}^{A}$ and $\hat{\pi}_{n}^{B}$ are approximated via standard particle filtering on the constrained models using $N$ particles, and we denote the Monte Carlo approximate filters as $\hat{\pi}_{n}^{A,N}$ and $\hat{\pi}_{n}^{B,N}$, respectively.
Figure 1 illustrates the exponential stability predicted by Theorem 3.1. Specifically, in Figure 1(a), we display the (approximate) total variation distances $\left\|\pi_{n}^{A,N_{1}}-\pi_{n}^{B,N_{1}}\right\|_{\rm tv}$ and $\left\|\pi_{n}^{A,N_{2}}-\pi_{n}^{B,N_{2}}\right\|_{\rm tv}$, obtained with $N_{1}=10,240$ and $N_{2}=1,024,000$ particles, averaged over 128 independent simulation runs. The prior distributions for this experiment are almost mutually singular, namely $\pi_{0}^{A}=\mathcal{N}(X_{0}-1,\dfrac{1}{4}I_{dx})$ and $\pi_{0}^{B}=\mathcal{N}(X_{0}+1,4I_{dx})$. The total variation distances are plotted versus the observation index (discrete time $n$). We see that the total variation distance between the constrained posterior measures with different priors decays exponentially fast, until it hits a Monte Carlo error floor, which is lower as we increase the number of particles in the PF.
@@
Figure 1: (a) The distances $\left\|\hat{\pi}_{n}^{A,N_{1}}-\hat{\pi}_{n}^{B,N_{1}}\right\|_{\rm tv}$ and $\left\|\hat{\pi}_{n}^{A,N_{2}}-\hat{\pi}_{n}^{B,N_{2}}\right\|_{\rm tv}$ versus observation index $n$, with number of particles $N_{1}=10,240$ and $N_{2}=1,024,000$ and barrier parameter $\varrho=2$. (b) The distance $\left\|\hat{\pi}_{n}^{A,N}-\hat{\pi}_{n}^{B,N}\right\|_{\rm tv}$ with $N=1,024,000$, barrier parameter $\varrho=2$ and the nonlinear fit $\frac{(1-\hat{\gamma}^{2})^{n}}{\hat{\gamma}^{2}}\left\|\hat{\pi}_{0}^{A}-\hat{\pi}_{0}^{B}\right\|_{\rm tv}$ of the contraction bound in (11).
Next, let us take a sequence of three constraints ${\mathcal{C}}^{l}$, $l=1,2,3$, which differ in the choice of the radius parameter $\varrho$ in the barrier function. In particular, we set $\kappa=1$ and $\mu=10$ for all three constraints, while $\{\varrho_{1},\varrho_{2},\varrho_{3}\}=\{0.5,~~2,~~4\}$. We then repeat the same experiment as in Figure 1, to obtain the decay of the total variation distance versus discrete time $n$.
The results are displayed in Figure 2(a). It illustrates how the volume of the constraint sets ${\mathcal{C}}_{n}^{l}$, determined by the barrier radius $\varrho_{l}$ affects the stability of the constrained filter. As $\varrho_{l}$ decreases, constrained sets become smaller, leading to a stronger contraction of the posterior measures and, hence, a faster decay of the total variation distance. The fitted decay rates are $\{\hat{\gamma}_{1},\hat{\gamma}_{2},\hat{\gamma}_{3}\}=\{0.518,0.505,0.487\}$, where larger $\gamma$ yields a faster decay.
@@ 5.4 Volume of constraint sets
In these experiments we study the effect the volume of the constraint sets on the accuracy of the barrier-constrained filters. Similar to the previous experiments we take stochastic Lorenz 96 model with state dimension $d_{x}=10$, $d_{y}=6$ observations at every time $t_{n}$, and noise scale parameters $\sigma_{x}^{2}=\frac{1}{2}$ and $\sigma_{y}^{2}=\frac{1}{4}$. The barrier method is implemented with parameters $\kappa=1$ and $\mu=10$, and radii $\{\varrho_{1},\varrho_{2},\varrho_{3},\varrho_{4}\}=\{0.5,1,2,4\}$. We note again that the radius $\varrho_{l}$ determines the volume of the sets ${\mathcal{C}}_{n}^{l}$. With this setup, we run a standard PF with $N=50,000$ particles for reference, to approximate an “optimal” posterior mean estimator, a standard PF with just $N=500$ particles, and several barrier-constrained PFs (for $\varrho_{1},\ldots,\varrho_{4}$) with $N=500$ particles.
Figure 2(b) displays the results in terms of the normalised mean square error (NMSE) between between the posterior mean estimate $\hat{x}_{n}^{N}=\frac{1}{N}\sum_{i=1}^{N}x_{n}^{i}$ output by each filter and the true state $X_{n}$, namely ${\rm NMSE}_{n}=\frac{T\|X_{n}-\hat{x}_{n}^{N}\|^{2}}{\sum_{m=1}^{T}\|X_{m}\|^{2}}$, where $T$ is the number of observation times. The standard PF with $N=500$ can be expected to lose track of the system state after less than 20 discrete time steps (for $n<20$) while the barrier-constrained filters perform close to the optimal posterior mean estimator for the complete simulation, with $50$ discrete time steps (the results are the same for longer simulations). The NMSE curves have been averaged over 128 independent simulation runs.
@@
Figure 2: (a) The (discretised) distance $\left\|\pi_{k}^{A,N}-\pi_{k}^{B,N}\right\|_{\rm tv}$ computed using standard PFs with $N=10,240$ particles and barrier parameters $\{\varrho_{1},\varrho_{2},\varrho_{3}\}=\{0.5,2,4\}$. (b) NMSE versus observation index (discrete time $n$) for the standard PF (with $N=500$ and $N=50,00$ particles) and barrier-constrained PFs with different constraints (parametrised by the barrier radii $\{\varrho_{1},\varrho_{2},\varrho_{3},\varrho_{4}\}=\{0.5,1,2,4\}$.

### v2

@@ 4.1 State space model
Let us assume that the signal of interest is a $d_{x}$-dimensional Itô process $\mathtt{X}(t)$ in $(\Omega,{\mathcal{F}},\mathbb{P})$, described by the stochastic differential equation (SDE)
@@
with initial condition $X_{n-1}=x^{\prime}$. For example, a standard Euler-Maruyama scheme yields
@@
The time grid (17) corresponds to the time instants where observations are collected and a likelihood function $g_{n}$ is available. We assume that the observations have the form
@@ 4.2 Constrained kernels
Let ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ be a constraint as described in Section 3.1. Sampling the kernel $\hat{K}_{n}(x^{\prime},dx)$ in (7) is equivalent to simulating paths of the Itô process $\mathtt{X}(t)$ that evolves from $\mathtt{X}(t_{n-1})=x^{\prime}$ up to a point in the set ${\mathcal{C}}_{n}$ at time $t_{n}$, i.e., we simulate $\mathtt{X}(t_{n})$ conditional on the initial condition $\mathtt{X}(t_{n-1})=x^{\prime}$ and the terminal condition $\mathtt{X}(t_{n})\in{\mathcal{C}}_{n}$.
One straightforward approach is to simulate trajectories of the original SDE (16) and then reject the samples $\mathtt{X}(t_{n})\notin{\mathcal{C}}_{n}$. However this can be extremely inefficient as the dimension $d_{x}$ increases. An alternative approach is to “guide” the Itô process to generate samples contained in ${\mathcal{C}}_{n}$ at time $t_{n}$. It is well known that the Doob $h$-transform [Rogers00]
@@
(19) ${\sf d}\mathtt{X}^{n}(t)=\left[a(\mathtt{X}^{n}(t),t)+\sigma^{2}(\mathtt{X}^{n}(t),t)\nabla_{x}\log h(\mathtt{X}^{n}(t),t)\right]{\sf d}t+\sigma(\mathtt{X}^{n}(t),t){\sf d}\mathtt{W}(t),$
with $\mathtt{X}^{n}(t_{n-1})=x^{\prime}$, describes a modified Itô process $\mathtt{X}^{n}(t)$ for $t\in[t_{n-1},t_{n}]$, and the law of $\mathtt{X}^{n}(t)$ coincides with the law of $\mathtt{X}(t)$ conditional on $\mathtt{X}(t_{n-1})=x^{\prime}$ and $\mathtt{X}(t_{n})\in{\mathcal{C}}_{n}$. Unfortunately, $h(X,t)$ is intractable except for a few particular cases (and also expensive to approximate).
@@
be a flow of superlevel sets of the Doob $h$-transform generated by some function $r:[t_{n-1},t_{n}]\mapsto(0,\infty)$, such that $\lim_{t\to t_{n}}r(t)=\infty$ to ensure $S_{n}(t_{n})={\mathcal{C}}_{n}$. Given $r(t)$, there is some $\epsilon_{r}\geq 0$ such that
@@
Assuming $\epsilon_{r}<<1$, the deterministic flow $S_{n}(t)$ captures most of the probability mass of the constrained process $\mathtt{X}^{n}(t)$ in Eq. (19). Intuitively, we can interpret it as a “high probability tube” connecting the initial condition $\mathtt{X}(t_{n-1})=x^{\prime}$ with the terminal set ${\mathcal{C}}_{n}$.
@@ 4.3 Barrier-constrained kernels
While approximating the Doob $h$-transform $h(x,t)$ is possible (several recent papers [Chopin23, Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25] have tackled this computation, e.g., using local linearization techniques or neural networks), we explore a simpler alternative where we construct a modified SDE that mimics the structure of (19) but replaces $\log h(x,t)$ by a simple-to-evaluate, differentiable “barrier function” $\log b(x,t)$. This barrier is designed to
@@
(20) ${\sf d}\mathtt{X}^{n}(t)=\left[a(\mathtt{X}^{n}(t),t)-\mu\sigma^{2}(\mathtt{X}^{n}(t),t)\nabla_{x}\log b(\mathtt{X}^{n}(t),t)\right]{\sf d}t+\sigma(\mathtt{X}^{n}(t),t){\sf d}\mathtt{W}(t),$
where $\mathtt{X}^{n}(t_{n-1})=x^{\prime}$ and $\mu>0$ is a parameter that controls the strength of the penalization when $\mathtt{X}^{n}(t)\notin\hat{S}_{n}(t)$.
Let ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:\log g_{n}(x)\geq\log\zeta_{n}\}$ be a superlevel set of the likelihood $g_{n}$, in agreement with the assumptions of Theorem 3.6. This construction implies that the sets $\hat{S}_{n}(t)$ depend on the log-likelihood $g_{n}$, hence it is useful to make the barrier depend on this quantity as well. This is relatively simple and can be done in a number of ways. In this paper we use a shifted soft-plus function with parameters $\varrho$ and $\kappa\geq 1$,
@@
Note that, for large $\kappa$, $u_{\varrho,\kappa}(z)\approx 0$, when $z<\varrho$, and $u_{\varrho,\kappa}(z)\approx z-\varrho$, when $z>\varrho$. The barrier function is then constructed as
@@ 5.1 Stochastic Lorenz 96 model
where $U_{n}\sim{\mathcal{N}}(0,\sigma_{y}^{2}I_{d_{y}})$, $H$ is a $d_{x}\times d_{y}$ matrix and $X_{n}=\mathtt{X}(t_{n})$. For each independent simulation trial, we generate $H$ randomly as
@@ 5.2 Barrier function
where $Y_{n}=y_{n}$ is the observation at time $t_{n}=n\Delta_{o}$. The constraint consists of superlevel sets ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:\log g_{n}(x)\geq\log\zeta\}$, where $\zeta\geq 0$ is the level. The tube $\hat{S}_{n}(t)$ is constructed as
$S_{n}(t):=\{x\in{\mathcal{X}}:\log g^{n}(x,t)\geq\log\zeta\},\quad t\in[t_{n-1},t_{n}],$
@@
Then, by construction, $g^{n}(x,t_{n-1})=g_{n-1}(x)$, $g^{n}(x,t_{n})=g_{n}(x)$, $\hat{S}_{n}(t_{n-1})={\mathcal{C}}_{n-1}$ and $\hat{S}_{n}(t_{n})={\mathcal{C}}_{n}$.
@@
$\hat{S}_{n}(t)=\{x\in{\mathcal{X}}:\log g^{n}(x,t)\geq\log\zeta\}=\left\{x\in{\mathcal{X}}:\|y^{n}(t)-H^{\top}x\|\leq\sqrt{-\frac{\log\zeta}{2\sigma_{y}^{2}}}\right\},$
hence, $\hat{S}_{n}(t)$ is the preimage (on the state space $\mathcal{X}$}) of the ball $\{y\in{\mathcal{Y}}:\|y^{n}(t)-y\|\leq R\}$, with radius $R=\sqrt{-\frac{\log\zeta}{2\sigma_{y}^{2}}}$ (on the observation space $\mathcal{Y}$}). The barrier function in (21) becomes
@@
where $\varrho=-\log\zeta$, and
@@ 5.3 Stability of constrained filters
From the SSMs ${\mathcal{S}}^{A}$ and ${\mathcal{S}}^{B}$ we construct constrained models $\hat{\mathcal{S}}^{A}$ and $\hat{\mathcal{S}}^{B}$. The constrained kernels $\hat{K}_{n}$ are implemented using the barrier method of Section 4.3 with parameters $\varrho=2$, $\kappa=1$ and $\mu=10$. We simulate the Lorenz 96 model with dimension $d_{x}=10$, with $d_{y}=6$ observations at every time $t_{n}$, and noise scale parameters $\sigma_{x}^{2}=\frac{1}{2}$ and $\sigma_{y}^{2}=\frac{1}{4}$. The discretisation of the total variation distance is carried out with $r=6$ and $r_{\rm sub}=3$. The filters $\hat{\pi}_{n}^{A}$ and $\hat{\pi}_{n}^{B}$ are approximated via a standard PF on the constrained models using $N$ particles, and we denote the Monte Carlo approximate filters as $\hat{\pi}_{n}^{A,N}$ and $\hat{\pi}_{n}^{B,N}$, respectively.
Figure 1 illustrates the exponential stability predicted by Theorem 3.1. Specifically, in Figure 1(a), we display the (approximate) total variation distances $\left\|\pi_{n}^{A,N_{1}}-\pi_{n}^{B,N_{1}}\right\|_{\rm tv}$ and $\left\|\pi_{n}^{A,N_{2}}-\pi_{n}^{B,N_{2}}\right\|_{\rm tv}$, obtained with $N_{1}=10,240$ and $N_{2}=1,024,000$ particles, averaged over 128 independent simulation runs. The prior distributions for this experiment are almost mutually singular, namely $\pi_{0}^{A}=\mathcal{N}(X_{0}-1,\dfrac{1}{4}I_{dx})$ and $\pi_{0}^{B}=\mathcal{N}(X_{0}+1,4I_{dx})$. The total variation distances are plotted versus the observation index (discrete time $n$). We see that the distance between the constrained filters with different priors decays exponentially fast, until it hits a Monte Carlo error floor, which is lower as we increase the number of particles $N$.
@@
Figure 1: (a) The distances $\left\|\hat{\pi}_{n}^{A,N_{1}}-\hat{\pi}_{n}^{B,N_{1}}\right\|_{\rm tv}$ and $\left\|\hat{\pi}_{n}^{A,N_{2}}-\hat{\pi}_{n}^{B,N_{2}}\right\|_{\rm tv}$ versus observation index $n$, with number of particles $N_{1}=10,240$ and $N_{2}=1,024,000$ and barrier parameter $\varrho=2$. (b) The distance $\left\|\hat{\pi}_{n}^{A,N_{2}}-\hat{\pi}_{n}^{B,N_{2}}\right\|_{\rm tv}$, barrier parameter $\varrho=2$ and the nonlinear fit $\frac{(1-\hat{\gamma}^{2})^{n}}{\hat{\gamma}^{2}}\left\|\hat{\pi}_{0}^{A}-\hat{\pi}_{0}^{B}\right\|_{\rm tv}$, with $\hat{\gamma}=0.496$, of the contraction bound in (10).
Next, let us take a sequence of three constraints ${\mathcal{C}}^{l}$, $l=1,2,3$, which differ in the choice of the radius parameter $\varrho$ in the barrier function. In particular, we take $\{\varrho_{1},\varrho_{2},\varrho_{3}\}=\{0.5,~~2,~~4\}$ and fix $\kappa=1$ and $\mu=10$. We then repeat the same experiment as in Figure 1, to obtain the decay of the total variation distance versus discrete time $n$.
The results are displayed in Figure 2(a). It illustrates how the volume of the constraint sets ${\mathcal{C}}_{n}^{l}$, determined by the barrier radius $\varrho_{l}$ affects the stability of the constrained filter. As $\varrho_{l}$ decreases, the sets ${\mathcal{C}}_{n}^{l}$ become smaller, leading to a stronger contraction of the posterior measures and, hence, a faster decay of the total variation distance. The fitted decay rates are $\{\hat{\gamma}_{1},\hat{\gamma}_{2},\hat{\gamma}_{3}\}=\{0.518,0.505,0.487\}$, where larger $\gamma$ yields a faster decay.
@@ 5.4 Volume of constraint sets
In these experiments we study the effect the volume of the constraint sets on the accuracy of the barrier-constrained filters. Similar to the previous experiments we take a stochastic Lorenz 96 model with state dimension $d_{x}=10$, $d_{y}=6$ observations at every time $t_{n}$, and noise scale parameters $\sigma_{x}^{2}=\frac{1}{2}$ and $\sigma_{y}^{2}=\frac{1}{4}$. The barrier method is implemented with parameters $\kappa=1$ and $\mu=10$, and radii $\{\varrho_{1},\varrho_{2},\varrho_{3},\varrho_{4}\}=\{0.5,1,2,4\}$. We note again that the radius $\varrho_{l}$ determines the volume of the sets ${\mathcal{C}}_{n}^{l}$. With this setup, we run a standard PF with $N=50,000$ particles for reference, to approximate an “optimal” posterior mean estimator, a standard PF with just $N=500$ particles, and several barrier-constrained PFs (for $\varrho_{1},\ldots,\varrho_{4}$) with $N=500$ particles.
Figure 2(b) displays the results in terms of the normalized mean square error (NMSE) between between the posterior mean estimate $\hat{x}_{n}^{N}=\frac{1}{N}\sum_{i=1}^{N}x_{n}^{i}$ output by each filter and the true state $X_{n}$, namely ${\rm NMSE}_{n}=\frac{M\|X_{n}-\hat{x}_{n}^{N}\|^{2}}{\sum_{m=1}^{M}\|X_{m}\|^{2}}$, where $M$ is the number of observation times. The standard PF with $N=500$ can be expected to lose track of the system state after less than 20 discrete time steps (for $n<20$) while the barrier-constrained filters perform close to the optimal posterior mean estimator for the complete simulation, with $50$ discrete time steps (the results are the same for longer simulations). The NMSE curves have been averaged over 128 independent simulation runs.
@@
Figure 2: (a) The (discretised) distance $\left\|\pi_{k}^{A,N}-\pi_{k}^{B,N}\right\|_{\rm tv}$ computed using standard PFs with $N=10,240$ particles and barrier parameters $\{\varrho_{1},\varrho_{2},\varrho_{3}\}=\{0.5,2,4\}$. (b) NMSE versus observation index (discrete time $n$) for the standard PF (with $N=500$ and $N=50,00$ particles) and barrier-constrained PFs with different constraints parametrized by the barrier radii $\{\varrho_{1},\varrho_{2},\varrho_{3},\varrho_{4}\}=\{0.5,1,2,4\}$.